OpenStudy (anonymous):

if the radius of the cylinder is 661.9 inches cubed and the diameter is 7 inches what is the height of the cylinder to the nearest tenth of an inch?

OpenStudy (anonymous):

why is the radius in inches cubed?

OpenStudy (anonymous):

perhaps you mean the volume

OpenStudy (anonymous):

also how could the diameter be less then the radius?

jimthompson5910 (jim_thompson5910):

the diameter is 7 inches, so the radius is exactly half of that: 7/2 =3.5 inches

jimthompson5910 (jim_thompson5910):

assuming you meant to say the volume was 661.9 cubic inches, this would mean that V = pi*r^2*h 661.9 = pi*(3.5)^2*h 661.9 = pi*12.25*h 661.9/(pi*12.25) = h h = 661.9/(pi*12.25) h = ???

OpenStudy (anonymous):

maybe i typed it wrong, heres what it looks like and im suppose to find it to the nearest tenth of an inch..

jimthompson5910 (jim_thompson5910):

oh the radius is 7, not the diameter thanks for the pic

jimthompson5910 (jim_thompson5910):

V = pi*r^2*h 661.9 = pi*(7)^2*h 661.9 = pi*49*h 661.9/(pi*49) = h h = 661.9/(pi*49) h = ??

jimthompson5910 (jim_thompson5910):

essentially what's going on is that we're using the volume of a cylinder formula, which is V = pi*r^2*h, plugging in the given info and solving for h

OpenStudy (anonymous):

hmmm I think understand it more now, thank you!

jimthompson5910 (jim_thompson5910):

so what do you get for h?

OpenStudy (anonymous):

I still didn't get it.

jimthompson5910 (jim_thompson5910):

where are you stuck? which line of my steps above doesn't make sense?

OpenStudy (anonymous):

im trying to figure out what to plug in for the formula haha

jimthompson5910 (jim_thompson5910):

the given volume is 661.9 so V = 661.9

jimthompson5910 (jim_thompson5910):

the radius is 7, so r = 7

jimthompson5910 (jim_thompson5910):

that explains line 2 in my work, making sense so far?

OpenStudy (anonymous):

somewhat

jimthompson5910 (jim_thompson5910):

to get line 3, i squared 7 to get 49

OpenStudy (anonymous):

I see

jimthompson5910 (jim_thompson5910):

line 4, i divided both sides by pi*49 to isolate h

jimthompson5910 (jim_thompson5910):

and in the next line I just flipped the equation

jimthompson5910 (jim_thompson5910):

the last line is where I wanted you to fill it in (by evaluating with a calculator)

OpenStudy (anonymous):

oh I see now

jimthompson5910 (jim_thompson5910):

yeah sorry I should have been more clear lol

jimthompson5910 (jim_thompson5910):

what does 661.9/(pi*49) evaluate to?

OpenStudy (anonymous):

I don't think I did it right ugh

jimthompson5910 (jim_thompson5910):

that's ok, post what you got

jimthompson5910 (jim_thompson5910):

don't be afraid to get it wrong that's what practice is for

OpenStudy (anonymous):

wait wait wait...is the / for dividing?

jimthompson5910 (jim_thompson5910):

correct, ex: 10/5 = 10 divided by 5 = 2

OpenStudy (anonymous):

okay, hold on

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

10329.01274?

jimthompson5910 (jim_thompson5910):

the 49 should be in the denominator

OpenStudy (anonymous):

im sooo confused

jimthompson5910 (jim_thompson5910):

what kind of calculator are you using?

OpenStudy (anonymous):

scientific

jimthompson5910 (jim_thompson5910):

and you're typing in 661.9/(pi*49) exactly as you see it?

jimthompson5910 (jim_thompson5910):

an alternative is to use google as a calculator (if your calculator isn't cooperating) so you would type "661.9/(pi*49)" without quotes into google to get https://www.google.com/search?hl=&q=661.9%2F%28pi*49%29&sourceid=navclient-ff&rlz=1B3GGLL_enUS420US420&ie=UTF-8

jimthompson5910 (jim_thompson5910):

what number does google give you as a result?

OpenStudy (anonymous):

4.2997...Aand so on

jimthompson5910 (jim_thompson5910):

good, so that's the height round that to the nearest tenth to get 4.3 inches

OpenStudy (anonymous):

woo! thank you!!

jimthompson5910 (jim_thompson5910):

you're welcome